Complete the table below for the equation y=2x-1 and draw a graph.
b)Estimate the x-value snd corresponding to y=2
c)find the x intercept.
d)find the y intercept.
Answer #1
Y=2X-1 2=2X-1 2X=2+1 2X=3 X=3/2 X=1.5 ANSWER. PROOF 2=2*1.5-1 2=3-1 2=2
answered by: Gayle
Answer #2
Complete the table below for the equation y=2x-1 and draw a graph. b)Estimate the x-value snd corresponding to y=2 2= 2x-1; 2x = 3; x =3/2 ----------------------- c)find the x intercept. If y=0, 2x-1=0 ; x = 1/2 is the x-intercept ------------------------ d)find the y intercept. If x=0, y = -1 is the y-intercept =============== Cheers, Stan H.
answered by: Dab
Answer #3

Y=2X-1 2=2X-1 2X=2+1 2X=3 X=3/2 X=1.5 ANSWER. PROOF 2=2*1.5-1 2=3-1 2=2

source: google
answered by: k1ngjayden